OpenStudy (anonymous):

Use calculator to find limits: when x approaches 0 for Lim (1-cosx)/x ; when x approaches infinity for Lim x/square root (3x^2-4)

hartnn (hartnn):

do u know the basic limit identity of limit x->0 sin x/x = ???

hartnn (hartnn):

or the question is to just use calcy and not do it analytically??

OpenStudy (anonymous):

the question asks to use calculator and give 2 decimal places

hartnn (hartnn):

still,u need to know limit x->0 sin x/x = ? do u know the value of this limit?

OpenStudy (anonymous):

umm, I don't know

hartnn (hartnn):

ok,\[\lim_{x \rightarrow 0}\frac{\sin x}{x}=1\] we are going to use this formula. first multiply (1+cos x) in numerator and denominator and tell me what u get in numerator.,after simplification??

OpenStudy (anonymous):

I got sinx/(x+cosx^2)

hartnn (hartnn):

seems incorrect, there was (1-cos x) in the numerator,i multiplied (1+cos x) so(1-cos x)(1+cos x)=1-cos^2 x = sin^2 x in the numerator.... did u get this?? there was x in denominator,i multiplied (1+cos x) to just get x(1+cos x) in denominator...

OpenStudy (anonymous):

yeah, I get it right, but I miss typed the sin^2 x in the numerator here.

hartnn (hartnn):

but u also got denominator as (x+cos^2 x) but actually,it is x(1+cos x) anyways,going ahead,i will split sin^2 x as sin x * sin x and write the limit as \[\lim_{x \rightarrow 0}\frac{\sin x}{x}\frac{\sin x}{1+\cos x}\] did u get this step??this is most important step.........

hartnn (hartnn):

but u also got denominator as (x+cos^2 x) but actually,it is x(1+cos x) anyways,going ahead,i will split sin^2 x as sin x * sin x and write the limit as \[\lim_{x \rightarrow 0}\frac{\sin x}{x}\frac{\sin x}{1+\cos x}\] did u get this step??this is most important step.........

OpenStudy (anonymous):

yeah, I get this one

hartnn (hartnn):

so now i will split the limits like this : \[\lim_{x \rightarrow 0}\frac{\sin x}{x}\lim_{x \rightarrow 0}\frac{\sin x}{1+\cos x}\] we know the value of 1st limit will be 1 the value of 2nd limit is evaluated by just putting x=0,u know the value of sin 0 and cos 0 ??

OpenStudy (anonymous):

So sin0=0 and cos0=1?

hartnn (hartnn):

yup correct,so just substitute that and tell me what u get the value of 2nd limit??

OpenStudy (anonymous):

Is it 0?

hartnn (hartnn):

yes it is :) all this work to get the final answer as 0!! correct to 2 decimal place,its 0.00 :P but u got all the steps?ask if u still have doubts anywhere...

OpenStudy (anonymous):

haha, that's ironic... it's a heavy process, but I get it now.

hartnn (hartnn):

hint for 2nd question,divide x in numerator and denominator....

hartnn (hartnn):

ask if u have difficulty solving the 2nd.....